OpenStudy (caitlinnr14):

WILL GIV MEDAL! Given the function f(x) = 2(x + 6), find x if f(x) = 22. 56 5 38 17

OpenStudy (debbieg):

Just plug in f(x)=22 and solve to find x: 22 = 2(x + 6)

OpenStudy (caitlinnr14):

oh ok thank you!

OpenStudy (anonymous):

x=5 f(x)=22 then 2(x+6)=22 implies x+6=11 therefore x=5

OpenStudy (caitlinnr14):

@DebbieG

OpenStudy (debbieg):

You're welcome. :)

OpenStudy (caitlinnr14):

so the answer would be 5 @DebbieG

OpenStudy (debbieg):

Yes, good job. :)

OpenStudy (caitlinnr14):

thxx

OpenStudy (anonymous):

I was told how to do it like this : f(x)=2(x+6) f(22)=2(22+6) f(22)=2(28) f(22)=56 & It was right every time ..

OpenStudy (caitlinnr14):

wait im confused @DebbieG @Sceenie

OpenStudy (debbieg):

@sceenie, that's a different question. THAT is finding f(22), which is DIFFERENT from "find x if f(x) = 22". If you were told to do that for THIS problem, that's wrong. But if it's "find f(22)" then it's right, but it's different problem. f(22) is the value of f(x) when x=22. "x when f(x)=22" is, well, the value of X when F(X)=22. Two different things.